b)  Answer the following stoichiometry questions by referring to the equation below:


2        KClO3  ---->  2 KCl + 3 O2


i.        If 1.50 mol of KClO3 decomposes, what is the mass of O2 that will be produced?

ii.       If 80.0 grams of O2 was produced, how many moles of KClO3 are decomposed?

iii.    Find the mass of KClO3 needed if we need to produce 2.75 mol of KCl.

1
Expert's answer
2021-02-23T04:25:44-0500

From the equation;

i. Mass of "O_2" produced

"1.50mol KClO_3" "\times" "3mol O_2\over 32.0gO_2" "\times" "122.5gKClO_2\over2mol KClO_2" "=8.61gO_2"


ii. "x" mol "KClO_3=80.0gO_2" "\times" "1mol O_2\over 32.0gO_2" "\times" "2mol KClO_3\over 3mol O_2"

"=1.67mol KClO_3"

iii. "x" g "KClO_3" "=2.75mol KCl" "\times" "2mol KCl\over 122.5gKClO_3" "\times" "74.55gKCl\over 2mol KCl"

"=1.67gKClO_3"